the sum of four times a number and three is the same as the difference of twice the number and eleven. find the number.

Answer by mananth(12270)

let number be x
4 times a number and 3
= 4x+3
..
difference of twice the number and 11
= 2x-11
..
2x-11=4x+3
4x-2x=-11-3
2x=-14
divide by2
x=-7
te number is -7